please tell the equation of tangents to a parabola .........

Grade:12

1 Answers

Vinayak Gupta
12 years ago

The equations are :

gradient form : y = mx + a/m where the point of contact is (a/m2,2a/m).

parametric form: ty = x + at2  where t=1/m.

cartesian form: yy1 = 2a(x+x1) where x1y1 is the point of contact.

 

 

approve the answer if helpful.